OpenStudy (anonymous):

Let logp/n=5 and logm/n=9. What is the relationship between P and M? A. P = 10,000M B. M = 0.0001P C. M = 4P D. M = 10,000P

OpenStudy (anonymous):

is it log(p/n) or (log p)/n

OpenStudy (anonymous):

the first option

OpenStudy (anonymous):

the first option

OpenStudy (anonymous):

i think D

OpenStudy (anonymous):

x = log y means y = a^x (e^x if ln)

OpenStudy (anonymous):

Yeah, if log(p/n) = 5 then 10^5 = p/n

OpenStudy (anonymous):

And following that logic for the second equation will give you a relation between m and p